The width of a rectangular lot is 75% of its length. If the perimeter is 700 meters, then what is the width and length?
Answer by mananth(16949)

let the length be x
width will be 0.75x
Perimeter = 2L+2W
2x+2*0.75x=700
2x+1.5x=700
3.5x=700
x=700/3.5
x=200 meters the length
width = 175 meters